# Dolphin: A Programmable Framework for Scalable Neurosymbolic Learning

### v1

@@ 3.3 Dolphin Provenances
As discussed earlier, each symbol in a distribution is associated with a batch of one or more tags. The Dolphin primitives define how to manipulate certain tags. For instance, Equations (1) and (3) specify the tags to be conjuncted or disjuncted. We now define the semantics of these operations.
The goal of such operations is to approximate the probabilities of the symbols in the output distribution as accurately as possible. This is achieved by using a mathematical framework called provenance semirings (Green et al., 2007). Provenance semirings provide generalized algebraic structure to propagate tags when computing over tagged data. In the case of Dolphin distributions, we can view the tags as representing the probabilities, and the data as the distribution’s symbols.
Designing and implementing provenances can be challenging since they must be accurate enough to capture the semantics of the symbolic program, while at the same time being coarse enough to maintain computational feasibility. Furthermore, the provenances must be differentiable to enable end-to-end training for neurosymbolic tasks. While neurosymbolic frameworks like Scallop (Li et al., 2023) implement differentiable provenances, they are not designed to leverage hardware accelerations or batched optimizations due to the CPU-bound nature of their implementations. We thus design vectorized provenances in Dolphin that are differentiable and enable GPU computations.
We simplify the definition of provenances from Scallop as a 5-tuple: $(T,\mathbf{0},\mathbf{1},\otimes,\oplus)$. Here, $T$ is the tag space, $\otimes:T\times T\to T$ is the conjunction operator with identity $\mathbf{0}$, and $\oplus:T\times T\to T$ is the disjunction operator with identity $\mathbf{1}$. We then implement two differentiable provenances in Dolphin: Differentiable Add-Mult Probabilities (DAMP) and Differentiable Top-K Proofs (DTKP). Table 1 summarizes the operations of these provenances.
@@
DTKP-AM $[0,1]\cup\{\infty,-\infty\}$ $\mathbf{\hat{0}}_{ij}=\begin{cases}\infty&i=1\\ -\infty&i>1\end{cases}$ $\mathbf{\hat{1}}_{ij}=\infty$ $\text{top}_{k}(\text{cat}(t,t^{\prime}))$ $\text{top}_{k}([\min(|t_{i}|,|t^{\prime}_{j}|)\mid(t_{i},t^{\prime}_{j})\in t\times t^{\prime}])$
@@
Differentiable Add-Mult Probabilities. Differentiable Add-Mult Probabilities (DAMP) is a popular technique that uses the probability space as its tag space: $T=[0,1]$. Its conjunction operation $\otimes$ is defined as the addition of probabilities, clamped at $\mathbf{1}$, and its disjunction operation $\oplus$ is defined as the multiplication of probabilities. The main assumption underlying the DAMP operations is that the input Distributions are mutually exclusive and independent. This assumption allows DAMP to compute probabilities extremely efficiently, as the operations are simple and can be easily vectorized.
Differentiable Top-$k$ Proofs. Differentiable Top-$k$ Proofs (DTKP) (Huang et al., 2021) was proposed to overcome the shortcomings of DAMP. This provenance tracks a set of up to $k$ proofs for each symbol. Each proof, in turn, denotes the set of input symbols necessary to derive the output symbol. These proofs are then used to compute the probabilities of the output symbols.
In Scallop, DTKP tags are converted into probabilities via differentiable weighted model counting (WMC). This form of DTKP is computationally hard and is by nature difficult to vectorize due to the varying sizes of proof sets and the WMC procedure. We therefore design a vectorized version of DTKP, called DTKP-AM (DTKP with Add-Mult), that can be efficiently computed on GPUs.
We first define the structure of tags in DTKP-AM in a manner that conforms to the constraints of PyTorch tensors. Each tag $t$ for a symbol $s$ is a 2-dimensional tensor of shape $(k,|I|)$, where $k$ is the maximum number of proofs to be retained and $I$ is an ordered list of all input symbols (symbols that are present in the input Distributions). Each row $t_{i}$ of $t$ corresponds to one of the tag’s $k$ proofs. Each element $t_{ij}$ thus represents the probability of the $j$th input symbol in the $i$th proof:
@@
where $p_{j}$ is the probability of the $j$th input symbol. The probability of each proof is then computed by taking the product of the normalized elements of that proof:
$\Pr(t_{i})=\prod_{j}\text{norm}(t_{ij})\quad\text{where}\quad\text{norm}(t_{ij})=\begin{cases}1&t_{ij}=+\infty\\ 0&t_{ij}=-\infty\\ t_{ij}&\text{otherwise}\end{cases}$
We next define the operations of DTKP-AM in Table 1. The $\oplus$ operation is defined as the union of two tag tensors $t$ and $t^{\prime}$ while $\otimes$ is defined as the element-wise minimum of the absolute values of all possible combinations of proofs in $t$ and $t^{\prime}$. In each case, the $\text{top}_{k}$ operation retains only upto $k$ proofs with the highest probabilities.
These definitions thus allow us to take advantage of the benefits of the DTKP provenance while enabling efficient computation on GPUs. To calculate the probability of the entire tag, DTKP-AM adds the probabilities of the individual proofs and clamps it at 1.
@@ 3.4 Building the Dolphin Program
The programmer specifies the neurosymbolic task using a Python program $P$ that uses Dolphin’s programming interface to connect the neural components (e.g., neural networks) with the symbolic components and operations. We call $P$ the symbolic program. Because $P$ is a Python program and Dolphin interfaces with PyTorch, Dolphin supports any PyTorch-based neural network(s), most Python language features, and custom user-defined functions. This feature enables greater flexibility and expressiveness in neurosymbolic programs than existing frameworks.
In addition to $P$, the programmer provides one or more neural networks $M_{1},\ldots,M_{k}$, and a dataset $\mathcal{D}$. Given these inputs, Dolphin extracts the computation graph that encodes how the neural network outputs are transformed using symbolic operations to produce a final result $D_{\text{res}}$. All computations in Dolphin are expressed using distribution objects $D_{i}$. Each Dolphin primitive (e.g., Apply) takes one or more distribution objects as inputs and applies a transformation to produce another distribution object. Because each distribution object $D_{i}$ only contains vectors of tags (or probabilities), the entire computation graph (including the neural network(s)) can be ported to a GPU for faster execution. During training, Dolphin optimizes over the standard objective function:
$\phi(\theta)=\min_{\theta}\sum_{(x,y)\in\mathcal{D}}\mathcal{L}(\text{$P$}(M_{\theta}(x)),y)$ (4)
Here $\mathcal{L}$ is the loss function, such as binary cross entropy. While Dolphin allows $P$ to take multiple neural networks as inputs, we show only one neural network model $M$ here for simplicity.
@@ 4 Experiments
  • RQ1: Scalability. How does Dolphin scale to complex problems and large datasets compared to existing baselines?
  • RQ2: Accuracy. Do models written in Dolphin converge to SOTA accuracies?
@@ 4.1 Benchmarks
MNIST-SumN. The MNIST-SumN (or briefly, SumN) task from (De Smet et al., 2024) takes as inputs $N$ handwritten digits from the MNIST dataset and returns their sum. We consider three versions of this task: small ($N=5$), medium ($N=10$), and large ($N=15$). The perception model for this task is a standard CNN while the symbolic program sums the outputs of this model.
Hand-Written Formula (HWF). The HWF task from Li et al. (2020) takes as input a set of images of handwritten digits and arithmetic operators representing a formula. The task is to evaluate the formula and return the result. We consider three versions of HWF: small (formulas of length up to 7), medium (formulas of length up to 15), and large (formulas of length up to 19). Similar to MNIST-SumN, the perception model for this task is a standard CNN. The Dolphin program for this task builds strings of formulae from the outputs of the neural model and evaluates them using Python’s eval function, demonstrating the ability of Dolphin to support black-box functions.
PathFinder. PathFinder (or Path) from  Tay et al. (2020) tests the ability of an agent to reason over long-range dependencies within an input image. The image consisting of two dots and a sequence of curved and dashed lines. The task is to identify whether the two dots are connected via the lines. We consider three versions of this task based on the image size in pixels: small (32 x 32), medium (128 x 128), and large (256 x 256). The perception model for this task is also a standard CNN. However, the Dolphin program for this task is recursive, since it must search for paths between the two dots.
CLUTRR. In this task from Sinha et al. (2019), the input is a passage of text containing some information about several individuals and some of their relationships. The task is then to infer the relationship between two given individuals, which is not explicitly provided in the input. Hence, the Dolphin program for this task needs to reason over long chains of relationships recursively. We consider two versions of this task, where the training data contains relation chains of lengths up to 3 (small) or 4 (medium). We use Roberta-base (Liu, 2019) as the language model for this task.
Mugen. In this task from Hayes et al. (2022), the input is a video of gameplay footage that is 3.2 seconds long and a natural language passage captioning the video. The goal is to measure how aligned the text is with the video. This task therefore has two variants: Mugen-TVR, where the model must retrieve the video that best aligns with the text, and Mugen-VTR, where the model must retrieve the text that best aligns with the video. This benchmark tests the ability of the model to reason over multimodal data. We consider two versions of this task: small, with 1000 training samples, and medium, with 5000 training samples. We use a combination of DistilBert (Sanh, 2019) and S3D (De Smet et al., 2024) as the perception model for the text and video inputs respectively.
@@ 4.2 Experimental Setup and Baselines
Setup. All experiments, with the exception of the CLUTRR benchmark, were run on machines with two 20-core Intel Xeon Gold 6248 CPUs, four NVIDIA GeForce RTX 4352 GPUs, and 768 GB RAM. Since the CLUTRR benchmark requires a larger amount of GPU memory, it was run on a machine with 8 NVIDIA A100 40GB GPUs instead. We ran each tool thrice until convergence or until a timeout of 10 hours was reached and report the average best accuracy and training time. The code of Dolphin and the benchmarks is provided in the supplementary material.
Baselines. We select Scallop (Li et al., 2023), a contemporary state-of-the-art neurosymbolic framework supporting differentiable programming optimized to run on the CPU and use multiple cores to parallelize its computations. We also choose two sampling-based gradient approximation methods, ISED (Solko-Breslin et al., 2024) and IndeCateR+ (De Smet et al., 2024). We compare Dolphin against Scallop on all benchmarks, and against ISED and IndeCateR on MNIST-SumN and HWF.
@@ 4.3 RQ1: Scalability
Task Dolphin Scallop ISED IndeCateR+
$\text{T}_{\text{total}}$ $\text{T}_{\text{epoch}}$ $\text{T}_{\text{total}}$ $\text{T}_{\text{epoch}}$ $\alpha$ $\text{T}_{\text{total}}$ $\text{T}_{\text{epoch}}$ $\alpha$ $\text{T}_{\text{total}}$ $\text{T}_{\text{epoch}}$ $\alpha$
SumN (S) 78.33 15.67 923.78 184.76 11.80 299.63 59.93 3.82 3.16e3 632.52 40.36
SumN (M) 144.92 14.49 3.41e3 341.57 23.57 2.16e3 216.54 14.94 2.63e4 2.63e3 181.88
SumN (L) 220.47 14.70 7.41e3 493.87 33.60 9.8e3 653.39 44.45 TO 2.27e3 154.37
HWF (S) 3.17e3 158.79 9.99e3 499.57 3.15 1.58e4 790.04 4.97 1.76e4 877.92 5.52
HWF (M) 1.46e4 731.67 TO 1.49e4 20.41 TO 6.83e3 9.34 TO 3.6e3 4.93
HWF (L) 2.42e4 1.21e3 TO 3.92e5 323.21 TO 1.05e4 8.61 TO 5.59e3 4.61
Path (S) 1.08e4 1.08e3 2.2e4 2.2e3 2.03 N.A.
Path (M) 1.79e4 1.79e3 TO 4.17e3 2.32
Path (L) 1.94e4 1.94e3 TO 1.12e4 5.81
CLUTRR (S) 1.54e3 154.85 4.29e3 429.97 2.77 N.A.
CLUTRR (M) 2.91e3 291.36 7.83e3 783.11 2.69
Mugen (S) 3.62e3 180.80 1.34e4 133.68 0.74 N.A.
Mugen (M) 1.78e3 890.34 TO 634.86 0.71
Table 2 presents the training times for Dolphin and baselines on all benchmarks, along with the total training time ($T_{\textit{total}}$), time per epoch ($T_{\textit{epoch}}$), and the scaling factor $\alpha$. The scaling factor is the ratio of the per epoch times of the baselines to Dolphin. We observe that in almost all cases, Dolphin shows significant improvements in training times, training up to about 300x faster than Scallop, 44x faster than ISED, and 180x faster than IndeCateR+. On average, Dolphin reports a speedup of 36x times over all baselines. Further, Dolphin converges on all benchmarks, while the baselines time out on most of the Medium and Large versions of the benchmarks. These results thus demonstrate that, unlike existing tools, Dolphin can scale to complex problems and large datasets.
Dolphin trains slightly slower than Scallop on both versions of the Mugen benchmark. This is because the Dolphin program written for Mugen uses Python objects and operations that are not fully batchable across samples. In contrast, the Scallop program, which is written in a compiled and optimized language, runs around 1.3x faster than Dolphin on average per iteration. However, Dolphin requires only 20 epochs to converge whereas Scallop requires almost 1000 epochs (Li et al., 2023). As a result, Dolphin’s total training time is still significantly lower than Scallop’s (3̃.7x for Mugen(S)). We show the accuracy curves for Mugen in Appendix B.

### v2

@@ 2.2 Dolphin Provenances
The Dolphin primitives discussed above define how to conjunct or disjunct tags corresponding to the symbol manipulations, e.g. Equations (1) and (3). These tag operations are achieved by using a mathematical framework called provenance semirings (Green et al., 2007). Provenance semirings provide generalized algebraic structure to propagate probabilities over tagged data.
Designing and implementing provenances can be challenging since they must be accurate enough to capture the semantics of the symbolic program, while at the same time being coarse enough to maintain computational feasibility. Furthermore, the provenances must be differentiable.
While neurosymbolic frameworks like Scallop (Li et al., 2023) implement differentiable provenances, they are not designed to leverage hardware accelerations or batched optimizations due to the CPU-bound nature of their implementations. Frameworks like LTN use t-norms that are more amenable to vectorization, but lack support for more complex provenances such as Differentiable Top-$k$ Proofs (DTKP) (Huang et al., 2021). We thus design differentiable, vectorized provenances in Dolphin to enable GPU support.
We simplify the definition of provenances as a 5-tuple: $(T,\mathbf{0},\mathbf{1},\otimes,\oplus)$. Here, $T$ is the tag space, $\otimes:T\times T\to T$ is the conjunction operator with identity $\mathbf{0}$, and $\oplus:T\times T\to T$ is the disjunction operator with identity $\mathbf{1}$. We then implement two differentiable provenances in Dolphin: Differentiable Add-Mult Probabilities (DAMP) and Differentiable Top-K Proofs (DTKP). Table 1 summarizes the operations of these provenances. While building the neurosymbolic program, the developer may specify which provenance to use, satisfying the core principle of tunability.
@@
DTKP-AM $[0,1]\cup\{\infty,-\infty\}$ $\mathbf{\hat{0}}_{ij}=-\infty$ $\mathbf{\hat{1}}_{ij}=\begin{cases}\infty&i=1\\ -\infty&i>1\end{cases}$ $\text{top}_{k}(\text{cat}(t,t^{\prime}))$ $\text{top}_{k}([\min(|t_{i}|,|t^{\prime}_{j}|)\mid(t_{i},t^{\prime}_{j})\in t\times t^{\prime}])$
@@
Differentiable Add-Mult Probabilities. Differentiable Add-Mult Probabilities (DAMP) is a popular technique that uses the probability space as its tag space: $T=[0,1]$. Its conjunction operation $\otimes$ is defined as the product of probabilities, clamped at $\mathbf{1}$, and its disjunction operation $\oplus$ is defined as the sum of probabilities. The main assumption underlying the DAMP operations is that the input Distributions are mutually exclusive and independent. This assumption allows DAMP to compute probabilities extremely efficiently, as the operations are simple and can be easily vectorized.
Differentiable Top-$k$ Proofs. Differentiable Top-$k$ Proofs (DTKP) (Huang et al., 2021) was proposed to overcome the shortcomings of DAMP. This provenance tracks a set of up to $k$ proofs for each symbol. Each proof denotes the set of input symbols necessary to derive the output symbol. These proofs are then used to compute the probabilities of the output symbols. In Scallop, DTKP tags are converted into probabilities via differentiable weighted model counting (WMC). This form of DTKP, which we call DTKP-WMC, is computationally hard and is by nature difficult to vectorize due to the varying sizes of proof sets and the WMC procedure. We hence design a novel vectorized approximation of DTKP-WMC; we term DTKP-AM (DTKP with Add-Mult), that can be efficiently computed on the GPU.
We first define the structure of tags in DTKP-AM to conform to the constraints of PyTorch tensors. Each tag $t$ for a symbol $s$ is a 2-dimensional tensor of shape $(k,|I|)$, where $k$ is the maximum number of proofs to be retained and $I$ is an ordered list of all input symbols (symbols that are present in the input Distributions). Each row $t_{i}$ of $t$ corresponds to one of the tag’s $k$ proofs. Each element $t_{ij}$ thus represents the probability of the $j$th input symbol in the $i$th proof:
@@
where $p_{j}$ is the probability of the $j$th input symbol. The probability of each proof is then computed by taking the product of the normal:
$\Pr(t_{i})=\prod_{j}\text{norm}(t_{ij}),\ \text{where}\ \text{norm}(t_{ij})=\begin{cases}1&t_{ij}=+\infty\\ 0&t_{ij}=-\infty\\ t_{ij}&\text{otherwise}\end{cases}$
We next define the operations of DTKP-AM in Table 1. The $\oplus$ operation is defined as the union of two tag tensors $t$ and $t^{\prime}$ while $\otimes$ is defined as the element-wise minimum of the normalized elements of all possible combinations of proofs in $t$ and $t^{\prime}$. In each case, the $\text{top}_{k}$ operation retains only up to $k$ proofs with the highest probabilities.
These definitions thus allow us to take advantage of the benefits of the DTKP provenance while enabling efficient computation on the GPU. To calculate the probability of the entire tag, DTKP-AM adds the probabilities of the individual proofs and clamps it at 1. We provide a detailed discussion of DTKP-AM in Appendix A.
@@ 2.3 Building the Dolphin Program
The programmer specifies the neurosymbolic task using a Python program $P$, which integrates neural components with symbolic operations via Dolphin’s interface. Given a dataset $\mathcal{D}$ and one or more neural networks $M_{1},\ldots,M_{k}$, Dolphin constructs a computation graph where symbolic transformations occur on the CPU, and probabilistic computations, including neural network inference, are efficiently executed on the GPU. All computations leverage distribution objects $D_{i}$, enabling end-to-end differentiability and scalability. Training optimizes the objective function
$\phi(\theta)=\min_{\theta}\sum_{(x,y)\in\mathcal{D}}\mathbb{L}(\text{$P$}(M_{\theta}(x)),y),$
where $\mathbb{L}$ is the loss function (e.g., binary cross entropy).
@@ 3 Experiments
  • RQ1: Scalability. Can Dolphin scale to tasks and datasets beyond the scope of existing SOTA frameworks?
  • RQ2: Accuracy. Do models written in Dolphin converge to SOTA accuracies in less training time?
@@ 3.1 Benchmarks
We describe the benchmarks used to evaluate Dolphin and give additional information about the experiment setup and Dolphin code for each benchmark in Appendix D.
MNIST SumN. The MNIST SumN (or briefly, SumN) task from (De Smet et al., 2024) takes as inputs $N$ handwritten digits from the MNIST dataset and returns their sum. We consider three versions of this task: SumN-5 ($N=5$), SumN-10 ($N=10$), and SumN-15 ($N=15$).
Hand-Written Formula (HWF). The HWF task from (Li et al., 2020) takes as input a set of images of handwritten digits and arithmetic operators representing a formula. The task is to evaluate the formula and return the result. We consider three versions based on formula length: HWF-7 (up to 7), HWF-15 (up to 15), and HWF-19 (up to 19).
PathFinder. PathFinder (or Path) (Tay et al., 2021) tests the ability of an agent to reason over long-range dependencies within an image of two dots and a sequence of curved and dashed lines. The task is to identify whether the two dots are connected via the lines. We consider three versions based on the image size in pixels: Path-32 (32 x 32), Path-128 (128 x 128), and Path-256 (256 x 256).
CLUTRR. In this task from (Sinha et al., 2019), given some text containing information about several individuals and some of their relationships, the model must infer the relationship between two given individuals, which is not explicitly provided in the input. We consider two versions, where the training data contains relation chains of lengths up to 3 (CLUTRR-3) or 4 (CLUTRR-4).
Mugen. In this task from (Hayes et al., 2022), given a 3.2 second long video of gameplay footage and text captioning the video, the goal is to measure how aligned the text is with the video. There are two variants: Mugen-TVR, where the model retrieves the video that best aligns with the text, and Mugen-VTR, where the model retrieves the text that best aligns with the video. We consider two versions of this task: 1K and 5K comprising 1000 and 5000 training samples.
@@ 3.2 Experimental Setup and Baselines
Setup. All experiments, except CLUTRR, were run on machines with two 20-core Intel Xeon Gold 6248 CPUs, four NVIDIA GeForce RTX 2080 Ti (11 GB) GPUs, and 768 GB RAM. Since CLUTRR demands more GPU memory due to running the RoBERTa model with a standard batch size of 16, all tools were run with a NVIDIA A100 40GB GPU. We ran each tool thrice until convergence or until a timeout of 10 hours was reached and report the average best accuracy and training time.
Baselines. We select Scallop (Li et al., 2023), a contemporary state-of-the-art neurosymbolic framework supporting differentiable programming optimized to run on the CPU in parallel using multiple cores. We also choose two sampling-based gradient approximation methods, ISED (Solko-Breslin et al., 2024) and IndeCateR+ (De Smet et al., 2024). We also include Logic Tensor Networks (LTN) (Serafini & Garcez, 2016), which combines first-order logic with continuous optimization by compiling logical constraints into a computation graph on the GPU. We compare Dolphin against Scallop on all benchmarks, and against ISED and IndeCateR+ on SumN and HWF. We compare SumN with LTN, but were unable to write HWF in LTN (explained in Appendix H.1). Since LTN also runs out of memory for a simpler benchmark like SumN-10, we do not compare against it for the more complex benchmarks of Path, CLUTRR, and Mugen. Similarly, we do not evaluate ISED and IndeCateR+ on these benchmarks, as ISED already fails to scale for simpler tasks like SumN-10 and HWF-7, while IndeCateR+ does not support their recursive structures.
Task Dolphin Scallop
$T_{\text{total}}$ $T_{\text{total}}$ $\alpha$
SumN-5 53.86 923.78 17.15
SumN-10 104.91 3.42e3 32.56
SumN-15 157.05 7.41e3 47.18
HWF-7 2.45e3 9.99e3 4.08
HWF-15 9.78e3 TO -
HWF-19 1.63e4 TO -
Path-32 1.29e4 2.2e4 1.71
Path-128 1.67e4 TO -
Path-256 1.97e4 TO -
CLUTRR-3 807.12 4.29e3 5.32
CLUTRR-4 923.86 7.83e3 8.48
Mugen-1K 3.81e3 1.34e4 3.52
Mugen-5K 1.62e4 TO -
Table 2: Training times for Dolphin and Scallop on all benchmarks (10-hour timeout). The scaling factor $\alpha$ is the ratio of the total training times of Scallop to Dolphin.
Task Dolphin LTN ISED IndeCateR+
$T_{\text{total}}$ $T_{\text{total}}$ $\alpha$ $T_{\text{total}}$ $\alpha$ $T_{\text{total}}$ $\alpha$
SumN-5 53.86 92.54 1.72 299.63 5.56 416.78 7.74
SumN-10 104.91 OOM - 2.17e3 20.64 385.65 3.68
SumN-15 157.05 OOM - 9.8e3 62.41 548.28 3.49
HWF-7 2.45e3 N.A. 1.58e4 6.45 1.35e4 5.51
HWF-15 9.78e3 TO - TO -
HWF-19 1.63e4 TO - TO -
Table 3: Training times for Dolphin, LTN, ISED, and IndeCateR+ on SumN and HWF (10-hour timeout). $\alpha$ is the ratio of training times of the baselines to Dolphin. OOM occurred on an NVIDIA GeForce RTX 2080 Ti (11 GB).
@@ 3.3 RQ1: Scalability
Table 2 presents the total training times ($T_{\text{total}}$) till convergence in seconds for Dolphin and Scallop across all benchmarks, alongside the scaling factor $\alpha$ (the ratio of the total training times of the baselines to Dolphin). Table 3 shows the same for the remaining baselines over SumN and HWF. The results demonstrate that Dolphin advances the state-of-the-art in neurosymbolic learning by scaling to more complex problems, e.g., larger versions of HWF, Path, and Mugen, that are beyond the reach of the other baselines which time out within 10 hours. Even for the other benchmarks where baselines do not time out, Dolphin achieves a scalability factor of up to 47x, 62x, and 3.49x against Scallop, ISED, and IndeCateR+ with an average speed up of 13.95x across all baselines for all benchmarks.
When the baselines do timeout, we use the training times per epoch to calculate a scalability factor $\alpha_{\text{epoch}}$ (Table 5, Appendix D). We see that among cases where baselines timeout, $\alpha_{\text{epoch}}$ is up to 280x for HWF-19 and is 36.4x faster on average. This results in Dolphin effectively training for more epochs in less time compared with the other baselines, which also allows it to converge to higher accuracies, as we see in Figure 3. We expand on these results in the next RQ.